# Controlling Counterfactual Harm in Decision Support Systems Based on Prediction Sets

### v1

@@ 1 Introduction
The principle of “first, do no harm” holds profound significance in a variety of professions across multiple high-stakes domains. For example, in the field of medicine, doctors swear an oath to prioritize their patient’s well-being or, in the legal justice system, preserving the innocence of individuals is paramount. As a result, in all of these domains, rules and guidelines have been established to prevent decision makers—doctors or judges—from making decisions that harm individuals—patients or suspects. In recent years, it has been increasingly argued that a similar principle should apply to decision support systems using machine learning algorithms in high-stakes domains [1, 2, 3, 4]. [^1] The European Unions’ AI act [5] mentions the term “harm” more than $35$ times and points out that, its crucial role in the design of algorithmic systems must be defined carefully.
@@
In machine learning for decision support, one of the main focus has been classification tasks. Here, the most studied setting assumes the decision support system uses a classifier to predict the value of a (ground-truth) label of interest and a human expert uses the predicted value to update their own prediction [9, 10, 11, 12, 13]. Unfortunately, in this setting, it is yet unclear how to guarantee that the (average) accuracy of the predictions made by an expert who uses the system is higher than the accuracy of the predictions made by the expert and the classifier on their own, what is often referred to as human-AI complementarity [14, 15, 16, 17]. In this context, a recent line of work [18, 19] have argued, both theoretically and empirically, that an alternative setting may enable human-AI complementarity. In this alternative setting, the decision support system helps a human expert by providing a set of label predictions, namely a prediction set, and asking them to always predict a label value from the set. The key principle is that, by restricting human agency, good performance does not depend on the expert developing a good sense of when to predict a label from the prediction set.
@@
Further related work. Our work builds upon further related work on set-valued predictors, counterfactual harm and algorithmic triage.
The literature on set-valued predictors seeks to design predictors that, for each instance, output a set of label values, namely a prediction set [22]. Set-valued predictors have not been typically designed nor evaluated by their ability to help human experts make more accurate predictions [21, 23, 24, 25, 26, 27, 28]. Only very recently, an emerging line of work has shown that conformal predictors, a specific type of set-valued predictors, may help human experts make more accurate predictions [18, 19, 29, 30, 31]. Within this line of work, the work most closely related to ours is by Straitouri et al. [18, 19], which has introduced the setting, and counterfactual and interventional monotonicity properties we build upon. However, in our work, instead of focusing on maximizing the average accuracy of the predictions made by humans as they do, we focus on controlling counterfactual harm.
@@
The literature on learning under algorithmic triage seeks to develop classifiers that make predictions for a given fraction of the samples and leave the remaining ones to human experts, as instructed by a triage policy [32, 33, 34, 35, 36, 37]. In contrast, in our work, for each sample, a classifier is used to construct a prediction set and a human expert needs to predict a label value from the set. In this context, it is also worth noting that learning under algorithmic triage has been extended to reinforcement learning settings [38, 39, 40, 41].
@@ 2 Decision support systems based on prediction sets
We consider a multiclass classification task in which, for each task instance, a human expert has to predict the value of a ground-truth label $y\in\mathcal{Y}=\{1,\ldots,L\}$. Then, our goal is to design a decision support system $\mathcal{C}:\mathcal{X}\rightarrow 2^{\mathcal{Y}}$ that, given a set of features $x\in\mathcal{X}$, helps the expert by narrowing down the label values to a subset of them $\mathcal{C}(x)\subseteq\mathcal{Y}$, namely a prediction set. Here, we focus on a setting in which the system asks the expert to always predict a label value $\hat{y}$ from the prediction set $\mathcal{C}(x)$. Note that, by restricting the expert’s agency, good performance does not depend on the human expert developing a good sense of when to predict a label from the prediction set. Moreover, we assume that the set of features, the ground-truth label and the expert’s prediction are sampled from an unknown fixed distribution [^2] We denote random variables with capital letters and realizations of random variables with lowercase letters., i.e., $x,y\sim P(X,Y)$ and $\hat{y}\sim P(\hat{Y}\,|\,X,Y,C_{\lambda}(X))$.
@@ 3 Counterfactual harm of decision support systems
where $\Lambda$, $U$ and $V$ are exogenous random variables and $f_{X}$, $f_{Y}$, $f_{\mathcal{C}}$ and $f_{\hat{Y}}$ are given functions. The exogenous variables $\Lambda$, $U$ and $V$ characterize the user-specified threshold, the expert’s individual characteristics and the data generating process, respectively. The function $f_{\mathcal{C}}$ is directly defined by Eq. 1, i.e., $f_{\mathcal{C}}(\Lambda,X)=\mathcal{C}_{\Lambda}(X)$. Further, as argued elsewhere [20], we can always find a distribution for the exogenous variables $\Lambda$, $U$ and $V$ as well as a functional form for the functions $f_{X}$, $f_{Y}$ and $f_{\hat{Y}}$ such that the distributions of the features, the ground-truth label and the expert’s prediction introduced in Section 2 are given by the observational distribution entailed by the SCM $\mathcal{M}$. For ease of exposition, we assume that, under no interventions, the distribution of the exogenous variable $\Lambda$ is $P(\Lambda)=\mathbbm{1}\{\Lambda=1\}$ and thus human experts make predictions on their own. Figure 1 shows a visual representation of our SCM $\mathcal{M}$.
Building upon the above characterization, we are now ready to formalize the following notion of counterfactual harm, which essentially compares the accuracy of a factual prediction made by an expert on their own against the counterfactual prediction that the same expert would have made had they use a decision support system $\mathcal{C}_{\lambda}$:
@@
For any $x,y,\hat{y}\sim P^{\mathcal{M}}$, the counterfactual harm that a decision support system $\mathcal{C}_{\lambda}$ would have caused, if deployed, is given by
@@
At this point, without further assumptions, we cannot expect to find the set $\Lambda(\alpha)$ because counterfactual harm lies within level three in the “ladder of causation” [20]. However, in what follows, we will show that, under certain assumptions, the average counterfactual harm $H(\lambda)$ is (partially) identifiable, i.e., it can be estimated (bounded) using observational data.
@@ 4 Counterfactual harm under counterfactual and interventional monotonicity
Under counterfactual monotonicity, for any $x\in\mathcal{X}$ and $\lambda,\lambda^{\prime}\in[0,1]$ such that $Y\in\mathcal{C}_{\lambda}(x)\subseteq\mathcal{C}_{\lambda^{\prime}}(x)$, if an expert has succeeded at predicting $Y$ using $\mathcal{C}_{\lambda^{\prime}}$, they would have also succeeded had they used $\mathcal{C}_{\lambda}$ and, conversely, if they has failed at predicting $Y$ using $\mathcal{C}_{\lambda}$, they would have also failed had they used $\mathcal{C}_{\lambda^{\prime}}$, while holding “everything else fixed”. More formally, counterfactual monotonicity is defined as follows:
@@
In what follows, we first show that, under the counterfactual monotonicity assumption, counterfactual harm is identifiable (we provide all proofs in Appendix A):
@@
Importantly, note that, in the above proposition, the lower bound on the left hand side of Eq. 9 matches the average counterfactual harm under the counterfactual monotonicity assumption and thus, holding “everything else fixed”, the average counterfactual harm under interventional monotonicity is always greater or equal than the average counterfactual harm under counterfactual monotonicity. Moreover, further note that the inequality condition of the interventional monotonicity assumption compares interventional probabilities and thus we can experimentally verify it (see Appendix D), lending support to using the above proposition to bound average counterfactual harm in high-stakes applications.
@@ 5 Controlling counterfactual harm using conformal risk control
Then, we leverage the above theorem and the fact that, under counterfactual monotonicity, the counterfactual harm is nonincreasing with respect to $\lambda$, as shown in Lemma 3 in Appendix A.5, to recover the largest harm-controlling set $\Lambda(\alpha)$:
@@
Under the interventional monotonicity assumption, rather than directly controlling the counterfactual harm, we will control the upper bound given by Eq. 9. Consequently, rather than recovering the largest harm-controlling set $\Lambda(\alpha)$, we will recover a harm-controlling set $\Lambda^{\prime}(\alpha)\subseteq\Lambda(\alpha)$. However, since the expression inside the expectation of the upper bound is nonmonotone with respect to $\lambda$, we cannot directly use it to set the value of the loss $\ell$ in conformal risk control. That said, since the first term of the expression is nonincreasing and the second term is nondecreasing, as shown in Lemmas 3 and 4 in Appendix A.5, we can apply conformal risk control separately for each term. For the first term, we use Theorem 1 because the term matches the counterfactual harm under the counterfactual monotonicity assumption. For the second term, we prove the following theorem:
@@ 6 Experiments
In this section, we use data from two different human subject studies to: a) evaluate the average counterfactual harm caused by decision support systems based on prediction sets; b) validate the theoretical guarantees offered by our computational framework (i.e., Corollaries 1 and 2); c) investigate the trade-off between the average counterfactual harm caused by decision support systems based on prediction sets and the average accuracy achieved by human experts using these systems. In what follows, we assume that counterfactual monotonicity holds. In Appendix E, we conduct experiments where we relax this assumption and assume instead that interventional monotonicity holds. [^5] All experiments are executed on a MacBook Pro laptop with an M1 processor, 16 GB of main memory and macOS operating system. The source code is publicly available at https://github.com/Networks-Learning/controlling-counterfactual-harm-prediction-sets.
@@
For each stratum of images, we apply our framework to decision support systems $\mathcal{C}_{\lambda}$ with different pre-trained classifiers, namely VGG19 [44], DenseNet161 [45], GoogleNet [46], ResNet152 [47] and AlexNet [48] after 10 epochs of fine-tuning, as provided by Steyvers et al. [42]. [^6] All classifiers and images are publicly available at https://osf.io/2ntrf. To this end, we randomly split the images (and human predictions) into a calibration set ($10$%), which we use to find the harm-controlling sets $\Lambda(\alpha)$ by applying Corollary 1, and a test set ($90$%), which we use to estimate the average counterfactual harm $H(\lambda)$ caused by the decision support systems $\mathcal{C}_{\lambda}$ as well as the average accuracy $A(\lambda)$ of the predictions made by a human expert using $\mathcal{C}_{\lambda}$. Here, since the dataset only contains predictions made by human participants on their own, we use the mixture of multinomial logit models (MNLs) introduced by Straitouri et al. [18] to estimate the average accuracy $A(\lambda)$. Refer to Appendix B for more details regarding the mixture of MNLs.
@@
Results. Figure 2 shows the average accuracy $A(\lambda)$ achieved by a human participant using $\mathcal{C}_{\lambda}$, as predicted by the mixture of MNLs, against the average counterfactual harm $H(\lambda)$ caused by $\mathcal{C}_{\lambda}$ for $\alpha\in\{0.01,0.05\}$ on the stratum of images with $\omega=110$ from the ImageNet16H dataset. Refer to Appendix C for results on other strata. Here, each point corresponds to a different $\lambda$ value and its coloring indicates the empirical probability that $\lambda$ is included in the harm-controlling set $\Lambda(\alpha)$. The results show several interesting insights. First, we find that, as long as $\lambda<1$, the decision support systems $\mathcal{C}_{\lambda}$ always cause some amount of counterfactual harm [^8] For $\lambda=1$, the decision support system $\mathcal{C}_{\lambda}$ does not cause harm because the prediction sets always contain all label values and thus human participants always make predictions on their own.. This suggests that, while restricting human agency enables human-AI complementarity, it inevitably causes (some amount of) counterfactual harm. Second, we find that the sets of $\lambda$ values provided by our framework are typically harm-controlling, i.e., they do not include $\lambda$ values such that $H(\lambda)>\alpha$. However, the sets are often conservative and do not include all the harm-controlling $\lambda$ values. Third, we find that there is a trade-off between accuracy and counterfactual harm and this trade-off is qualitatively consistent across decision support systems using different pre-trained classifiers. In fact, for $\alpha=0.01$, the decision support systems $\mathcal{C}_{\lambda}$ offering the greatest average accuracy $A(\lambda)$ are not harm-controlling.

### v2

@@ 1 Introduction
The principle of “first, do no harm” holds profound significance in a variety of professions across multiple high-stakes domains. For example, in the field of medicine, doctors swear an oath to prioritize their patient’s well-being or, in the legal justice system, preserving the innocence of individuals is paramount. As a result, in all of these domains, rules and guidelines have been established to prevent decision makers—doctors or judges—from making decisions that harm individuals—patients or suspects. In recent years, it has been increasingly argued that a similar principle should apply to decision support systems using machine learning algorithms in high-stakes domains [1, 2, 3, 4]. [^1] The European Unions’ AI act mentions the term “harm” more than $35$ times and points out that, its crucial role in the design of algorithmic systems must be defined carefully.
@@
In machine learning for decision support, one of the main focus has been classification tasks. Here, the most studied setting assumes the decision support system uses a classifier to predict the value of a (ground-truth) label of interest and a human expert uses the predicted value to update their own prediction [8, 9, 10, 11, 12]. Unfortunately, in this setting, it is yet unclear how to guarantee that the (average) accuracy of the predictions made by an expert who uses the system is higher than the accuracy of the predictions made by the expert and the classifier on their own, what is often referred to as human-AI complementarity [13, 14, 15, 16]. In this context, a recent line of work [17, 18] have argued, both theoretically and empirically, that an alternative setting may enable human-AI complementarity. In this alternative setting, the decision support system helps a human expert by providing a set of label predictions, namely a prediction set, and asking them to always predict a label value from the set. The key principle is that, by restricting human agency, good performance does not depend on the expert developing a good sense of when to predict a label from the prediction set. In this context, it is also worth noting that Google has recently developed a tool that uses patient history and skin condition images to provide decision support using prediction sets [19], and a study by Jain et al. [20] has found that physicians and nurses using this tool improved diagnoses for $1$ in every $8$ to $10$ cases.
@@
Further related work. Our work builds upon further related work on set-valued predictors, critiques of prediction optimization, counterfactual harm and algorithmic triage.
Set-valued predictors output a set of label values, namely a prediction set, rather than single labels [23]. However, set-valued predictors have not been typically designed nor evaluated by their ability to help human experts make more accurate predictions [24, 25, 26, 27, 28, 29]. Only very recently, an emerging line of work has shown that conformal predictors, a specific type of set-valued predictors, may help human experts make more accurate predictions [17, 18, 30, 31, 32, 33]. Within this line of work, the work most closely related to ours is by Straitouri et al. [17, 18], which has introduced the setting, and counterfactual and interventional monotonicity properties we build upon.
Prediction optimization has been recently put into question in the context of decision support [34, 35, 36]. More specifically, it has been argued that optimizing decision support systems to improve prediction accuracy does not always translate to better decision-making. Our work aligns with this critique since we argue that improving prediction accuracy may come at the cost of counterfactual harm.
@@
Learning under algorithmic triage seeks to develop classifiers that make predictions for a given fraction of the samples and leave the remaining ones to human experts, as instructed by a triage policy [37, 38, 39, 40, 41, 42]. In contrast, in our work, for each sample, a classifier is used to construct a prediction set and a human expert needs to predict a label value from the set. In this context, it is also worth noting that learning under algorithmic triage has been extended to reinforcement learning settings [43, 44, 45, 46].
@@ 2 Decision support systems based on prediction sets
We consider a multiclass classification task in which, for each task instance, a human expert has to predict the value of a ground-truth label $y\in\mathcal{Y}=\{1,\ldots,L\}$. Then, our goal is to design a decision support system $\mathcal{C}:\mathcal{X}\rightarrow 2^{\mathcal{Y}}$ that, given a set of features $x\in\mathcal{X}$, helps the expert by narrowing down the label values to a subset of them $\mathcal{C}(x)\subseteq\mathcal{Y}$, namely a prediction set. Here, we focus on a setting in which the system asks the expert to always predict a label value $\hat{y}$ from the prediction set $\mathcal{C}(x)$. Note that, by restricting the expert’s agency, good performance does not depend on the human expert developing a good sense of when to predict a label from the prediction set. Moreover, we assume that the set of features, the ground-truth label and the expert’s prediction are sampled from an unknown fixed distribution [^2] We denote random variables with capital letters and realizations of random variables with lowercase letters., i.e., $x,y\sim P(X,Y)$ and $\hat{y}\sim P(\hat{Y}\,|\,X,Y,\mathcal{C}(X))$.
@@
Remark. We would like to clarify that, if one sets the value of the threshold $\lambda$ to be roughly the $1-\alpha$ quantile of the empirical distribution of the scores $1-m_{y}(x)$ in a calibration set, then, the set valued predictor defined by Eq. 1 is equivalent to a vanilla conformal predictor with nonconformity scores $1-m_{y}(x)$ and coverage $1-\alpha$. Under this view, it becomes apparent that, by searching for $\lambda$ values in $[0,1]$ that are harm controlling, we are essentially searching for vanilla conformal predictors that are harm controlling. In this context, we would like to further clarify that our framework is agnostic to the choice of nonconformity score or more generally, the set-valued predictor, used to construct the prediction sets [47, 48, 49]. Motivated by this observation, in Appendix E, we include additional experiments where we evaluate our framework using a more complex set-valued predictor [49].
@@ 3 Counterfactual harm of decision support systems
where $\Lambda$, $U$ and $V$ are exogenous random variables and $f_{X}$, $f_{Y}$, $f_{\mathcal{C}}$ and $f_{\hat{Y}}$ are given functions. [^4] The functions $f_{X}$, $f_{Y}$, $f_{\mathcal{C}}$ and $f_{\hat{Y}}$ are causal mechanisms and not equations that can be manipulated [50]. The exogenous variables $\Lambda$, $U$ and $V$ characterize the user-specified threshold, the expert’s individual characteristics and the data generating process, respectively. The function $f_{\mathcal{C}}$ is directly defined by Eq. 1, i.e., $f_{\mathcal{C}}(\Lambda,X)=\mathcal{C}_{\Lambda}(X)$. Further, as argued elsewhere [21], we can always find a distribution for the exogenous variables $\Lambda$, $U$ and $V$ as well as a functional form for the functions $f_{X}$, $f_{Y}$ and $f_{\hat{Y}}$ such that the distributions of the features, the ground-truth label and the expert’s prediction introduced in Section 2 are given by the observational distribution entailed by the SCM $\mathcal{M}$. For ease of exposition, we assume that, under no interventions, the distribution of the exogenous variable $\Lambda$ is $P(\Lambda)=\mathbbm{1}\{\Lambda=1\}$ and thus human experts make predictions on their own. Figure 1 shows a visual representation of our SCM $\mathcal{M}$.
Building upon the above characterization, we are now ready to formalize the following notion of counterfactual harm, which essentially compares the accuracy of a factual prediction made by an expert on their own against the counterfactual prediction that the same expert would have made had they used a decision support system $\mathcal{C}_{\lambda}$:
@@
For any $x,y,\hat{y}\sim P^{\mathcal{M}}$, the counterfactual harm that a decision support system $\mathcal{C}_{\lambda}$ would have caused, if deployed, is given by [^5] To denote interventions in a counterfactual distribution, we follow the notation by Peters et al. [51]. Refer to Appendix A for a comparison to Pearl’s notation [50].
@@
At this point, we cannot expect to find the set $\Lambda(\alpha)$ because counterfactual harm lies within level three in the “ladder of causation” [21] and thus it is not identifiable from observational data without further assumptions. However, in what follows, we will show that, under certain assumptions, the average counterfactual harm $H(\lambda)$ is (partially) identifiable, i.e., it can be estimated (bounded) using observational data.
Comparison to Richens’s definition of counterfactual harm. Richens et al. [1] define counterfactual harm as follows:
$h_{\lambda}(x,y,\hat{y})=E_{\hat{Y}\sim P^{\mathcal{M}|\hat{Y}=\hat{y},X=x,Y=y}(\hat{Y})}[\max\{0,\mathbbm{1}\{\hat{Y}=y\}-\mathbbm{1}\{\hat{y}=y\}\}],$
where $x,y,\hat{y}\sim P^{\mathcal{M};do(\Lambda=\lambda)}$ and $\lambda=1$ is considered to be the default action in the language of Richens et al. [1]. This definition implicitly assumes that the system $\mathcal{C}_{\lambda}$ is deployed and it compares the factual prediction $\hat{y}$ made by an expert using the deployed system against the counterfactual prediction $\hat{Y}$ had the expert made on their own. On the contrary, our definition does not assume that the system $\mathcal{C}_{\lambda}$ is deployed and instead it compares the factual prediction $\hat{y}$ made by an expert on their own against the counterfactual prediction $\hat{Y}$ had the expert made using the system $\mathcal{C}_{\lambda}$.
@@ 4 Counterfactual harm under counterfactual and interventional monotonicity
Under counterfactual monotonicity, for any $x\in\mathcal{X}$ and $\lambda,\lambda^{\prime}\in[0,1]$ such that $Y\in\mathcal{C}_{\lambda}(x)\subseteq\mathcal{C}_{\lambda^{\prime}}(x)$, if an expert has succeeded at predicting $Y$ using $\mathcal{C}_{\lambda^{\prime}}$, they would have also succeeded had they used $\mathcal{C}_{\lambda}$ and, conversely, if they have failed at predicting $Y$ using $\mathcal{C}_{\lambda}$, they would have also failed had they used $\mathcal{C}_{\lambda^{\prime}}$, while holding “everything else fixed”. More formally, counterfactual monotonicity is defined as follows:
@@
In what follows, we first show that, under the counterfactual monotonicity assumption, counterfactual harm is identifiable (we provide all proofs in Appendix B):
@@
Importantly, note that, in the above proposition, the lower bound on the left hand side of Eq. 9 matches the average counterfactual harm under the counterfactual monotonicity assumption and thus, holding “everything else fixed”, the average counterfactual harm under interventional monotonicity is always greater or equal than the average counterfactual harm under counterfactual monotonicity. Moreover, further note that the inequality condition of the interventional monotonicity assumption compares interventional probabilities and thus we can experimentally verify it (see Appendix F), lending support to using the above proposition to bound average counterfactual harm in high-stakes applications.
@@ 5 Controlling counterfactual harm using conformal risk control
Then, we leverage the above theorem and the fact that, under counterfactual monotonicity, the counterfactual harm is nonincreasing with respect to $\lambda$, as shown in Lemma 3 in Appendix B.5, to recover the largest harm-controlling set $\Lambda(\alpha)$:
@@
Under the interventional monotonicity assumption, rather than directly controlling the counterfactual harm, we will control the upper bound given by Eq. 9. Consequently, rather than recovering the largest harm-controlling set $\Lambda(\alpha)$, we will recover a harm-controlling set $\Lambda^{\prime}(\alpha)\subseteq\Lambda(\alpha)$. However, since the expression inside the expectation of the upper bound is nonmonotone with respect to $\lambda$, we cannot directly use it to set the value of the loss $\ell$ in conformal risk control. That said, since the first term of the expression is nonincreasing and the second term is nondecreasing, as shown in Lemmas 3 and 4 in Appendix B.5, we can apply conformal risk control separately for each term. For the first term, we use Theorem 1 because the term matches the counterfactual harm under the counterfactual monotonicity assumption. For the second term, we prove the following theorem:
@@ 6 Experiments
In this section, we use data from two different human subject studies to: a) evaluate the average counterfactual harm caused by decision support systems based on prediction sets; b) validate the theoretical guarantees offered by our computational framework (i.e., Corollaries 1 and 2); c) investigate the trade-off between the average counterfactual harm caused by decision support systems based on prediction sets and the average accuracy achieved by human experts using these systems. In what follows, we assume that counterfactual monotonicity holds. In Appendix G, we conduct experiments where we relax this assumption and assume instead that interventional monotonicity holds. [^7] All experiments ran on a Mac OS machine with an M1 processor and 16GB Memory. An open-source implementation of our methodology is publicly available at https://github.com/Networks-Learning/controlling-counterfactual-harm-prediction-sets.
@@
For each stratum of images, we apply our framework to decision support systems $\mathcal{C}_{\lambda}$ with different pre-trained classifiers, namely VGG19 [54], DenseNet161 [55], GoogleNet [56], ResNet152 [57] and AlexNet [58] after 10 epochs of fine-tuning, as provided by Steyvers et al. [52]. [^8] All classifiers and images are publicly available at https://osf.io/2ntrf. To this end, we randomly split the images (and human predictions) into a calibration set ($10$%), which we use to find the harm-controlling sets $\Lambda(\alpha)$ by applying Corollary 1, and a test set ($90$%), which we use to estimate the average counterfactual harm $H(\lambda)$ caused by the decision support systems $\mathcal{C}_{\lambda}$ as well as the average accuracy $A(\lambda)$ of the predictions made by a human expert using $\mathcal{C}_{\lambda}$. Here, since the dataset only contains predictions made by human participants on their own, we use the mixture of multinomial logit models (MNLs) introduced by Straitouri et al. [17] to estimate the average accuracy $A(\lambda)$. Refer to Appendix C for more details regarding the mixture of MNLs and to Appendix D for additional experiments studying the relationship between average counterfactual harm $H(\lambda)$, average prediction set size and empirical coverage [^9] By empirical coverage, we refer to the fraction of the test samples for which the prediction sets include the ground truth label value..
@@
Results. Figure 2 shows the average accuracy $A(\lambda)$ achieved by a human participant using $\mathcal{C}_{\lambda}$, as predicted by the mixture of MNLs, against the average counterfactual harm $H(\lambda)$ caused by $\mathcal{C}_{\lambda}$ for $\alpha\in\{0.01,0.05\}$ on the stratum of images with $\omega=110$ from the ImageNet16H dataset. Refer to Appendix D for results on other strata. Here, each point corresponds to a different $\lambda$ value and its coloring indicates the empirical probability that $\lambda$ is included in the harm-controlling set $\Lambda(\alpha)$. The results show several interesting insights. First, we find that, as long as $\lambda<1$, the decision support systems $\mathcal{C}_{\lambda}$ always cause some amount of counterfactual harm [^11] For $\lambda=1$, the decision support system $\mathcal{C}_{\lambda}$ does not cause harm because the prediction sets always contain all label values and thus human participants always make predictions on their own.. This suggests that, while restricting human agency enables human-AI complementarity, it inevitably causes (some amount of) counterfactual harm. Second, we find that the sets of $\lambda$ values provided by our framework are typically harm-controlling, i.e., they do not include $\lambda$ values such that $H(\lambda)>\alpha$. However, the sets are often conservative and do not include all the harm-controlling $\lambda$ values due to estimation error in the the empirical estimate $\hat{H}_{n}(\lambda)$ of the average counterfactual harm using data from the calibration set. In Appendix D, we show that using larger calibration sets reduces the above mentioned estimation error and results in sets of $\lambda$ values that are less conservative. Third, we find that there is a trade-off between accuracy and counterfactual harm and this trade-off is qualitatively consistent across decision support systems using different pre-trained classifiers. In fact, for $\alpha=0.01$, the decision support systems $\mathcal{C}_{\lambda^{*}}$ offering the greatest average accuracy $A(\lambda^{*})$ are not harm-controlling.